Working together, Rick and Juanita can complete a job in 6 hours. It would take Rick 9 hours longer than Juanita to do the job alone. How long would it take Juanita alone?
Answer by checkley71(8403)

6=x*(x+9)/(x+x+9)
6=(x^2+9x)/(2x+9) cross multiply
6(2x+9)=x^2+9x
12x+54=x^2+9x
x^2+9x-12x-54=0
x^2-3x-54=0
(x-9)(x+6)=0
x-9 =0
x=9 answer.
proof
6=(9*18)/9+9+9
6=162/27
6=6